How to solve x^2+4kx-k by completing the square

Found 2 solutions by stanbon, rothauserc:
Answer by stanbon(75887)

solve x^2+4kx-k by completing the square
---
You cannot "solve" if you do not have a equation.
-----
I'll assume you meant x^2 + 4kx -k = 0
------------------------
x^2 + 4kx = k
x^2 + 4kx + (2k)^2 = k + (2k)^2
----
(x+2k)^2 = 4k^2+k
-----
x+2k = +/-sqrt(4k^2+k)
x = -2k +/-sqrt(4k^2+k)
----------------------------
Cheers,
Stan H.
---------------

Answer by rothauserc(4718)

x^2 +4kx -k = 0
x^2 +4kx = k
x^2 +4kx +4k^2 = k +4k^2
(x +2k)^2 = k +4k^2
x+2k = sqrt(k+4k^2)
x+2k = sqrt(k(1+4k))
x+2k = sqrt(k) * sqrt(1+4k)
x = sqrt(k) * sqrt(1+4k) - 2k